Please someone I have a big test and I am totally confused here
solve by the substitution method
7x+9y=35
x=10-2y

Answer by mananth(16949)

7x+9y=35
x=10-2y
substitute x in the first equation
7(10-2y) +9y = 35
70-14y+9y=35
-5y=-35
/-5
y=7
plug y in second equation
x=10-2*7
x=10-14
x=-4
(-4,7)